# Synthetic Biology Meets Neuromorphic Computing: Towards a Bio-Inspired Olfactory Perception System

### v1

@@ 3.2 Surface stabilization of Orco-OR complexes in synthetic cells
To address the challenge of the necessary co-design approach, we developed a quantitative model of the proposed assembly of lipid-membrane-stabilized Orco-OR ion channels on an electrode. In the model, a fixed number of ion channels generate a current that flows into the electrode impedance. The ion channel was modeled as a resistor with two conductance states, and the electrodes were considered as a parallel resistor-capacitor circuit. The cleft between the synthetic cell and the electrode acts as a resistive element and is relatively large due to the nanoscopic dimensions of the cleft (see Materials and Methods). Analysis of the model showed that the vesicle radius r is an important design parameter. The total charge in a synthetic cell scales with $r^{3}$, while the number of ion channels scales with $r^{2}$. Thus, we expect two regimes with different vesicle sizes: one regime dominated by the finite number of charges encapsulated in smaller vesicles, whereas signals from larger vesicles are sensitive to the densities of inserted ion channels. In addition, the overall coverage of signal-generating synthetic cells depends on the ratio of electrode size to vesicle size. In experiments, the sizes of vesicles are readily tuned between 100 nm and 100 $\mu$m, providing an opportunity to test the different regimes predicted by the model.
@@ 3.3 Selection methodology for odor receptors (OR) and synthetic dataset
To this end, we make use of a synthetic dataset based on measurements of odorant responses in Drosophila melanogaster. In the DoOR dataset, all recordings have been merged and normalized across studies, providing response data for 78 types of sensory units to 692 odorants. As we aim to reconstitute OR, we exclude units from the dataset which can not be clearly linked to a single OR type, such as responses from multiple receptor types or ionotropic receptors. This leaves our analysis with a total of 52 possible OR types.
We first need to restrict the number of ORs and odorants. The task is to find the minimal number of ORs we need to reconstitute in order to reliably differentiate between a given set of odors. This can be seen as a separation task in the space of responses of all receptors, see Figure 4 (left). We achieve this by calculating the ‘odor separability’, i.e. the Euclidean distance between the response vectors to the odor, and iteratively excluding the OR which contributes least to separability; see Section 5.4 for a detailed description of the method.
Our analysis yields the odor separability for different sets of odorants vs. the number of enabled ORs, as shown in Figure 4 (center). For illustration, we have chosen three sets of odors with differing degrees of relationship in terms of chemical structure (blue, orange, and green sets). As expected, all odor sets show better separability with larger number of different receptors; however, relative separability naturally depends on the combination of odors we are probing (note that results are normalized across odorant sets). Blue: three odorants from different functional groups (respectively: alcohols, esters, and lactones) show large relative separability. Orange: separability is reduced when considering less diverse functional groups (here: alcohols and aromatic alcohols). Green: odorants from a single functional group with similar OR responses are naturally less separable. Note that we control against the possibility of no ligand binding at all for a given odorant by including a ‘none’ odor; see Section 5.4 in Materials and Methods.
Most importantly, our algorithm tells us exactly which ORs need to be reconstituted for optimal separability of odors and subsequent discrimination of odors by our sensing platform. As an example, Figure 4 (right) shows the most relevant ORs for $N_{\text{OR}}=3$ for the aforementioned odor sets.
@@
Figure 4: Left: OR/odor space, exemplified by three smells that occupy a space defined by the responses of different ORs. Our algorithm selects OR for maximum separability in the OR space. Center: Normalized odor separability vs. number of Dm ORs for different sets of odorants. Right: ORs showing highest separability for each odor set for $N_{\text{OR}}=3$.
@@
For a single OR, the response to the binding of an odor molecule is modeled as a stochastic process with $N_{\text{ORCO}}$ ion channels. If the ligand binds to a closed Orco channel, the channel opens, and current flows out of the vesicle until stochastic unbinding occurs. We thus define a Markov model with the opening probability of Orco $p_{\text{open}}$ and closing with $p_{\text{close}}$. The open and closed states are converted to currents by comparison with the electrophysiological characterization of VUAA1-induced Orco currents; see Section 5.5 and Butterwick et al. 2018 [67], in particular Extended Data Fig. 2. Similarly, ion channel noise and low-pass filter properties of the vesicle membrane are modeled on the basis of electrophysiological characterization. To obtain the current flow of an OR of a single receptor type, we then sum over all of its Orco currents. For the electrode response, we assume that the electrode can probe individual ORs; thus, the sensor response to an odor molecule is given by a low-pass filter of the individual OR current, which we parametrize with the electrode model described in Section 3.2. From this procedure, we obtain a time series of voltages with sensor responses per odor–OR type combination, see Figure 5. The model is flexible with respect to which odorants and ORs to simulate, supporting the full DoOR 2.0.1 dataset of 692 odors and 52 OR types. In order to increase realism w.r.t. actual measurement conditions, the onset of odor presentation is chosen at random within the full simulation time window.
Figure 5: Simulated electrode responses to three odorants. Each trace shows the response of one synthetic cell comprising 10 Orco-OR complexes of the same receptor type; different receptor types are color-coded. We show the 7 most relevant OR types, as determined by our OR reduction algorithm. Opacity of traces is scaled with maximum voltage for better readability. Note that the odor is only presented during a portion of the overall time frame (presentation onset after 0.050.050.05\,0.05s, 0.80.80.8\,0.8s, 0.40.40.4\,0.4s, from top to bottom).
@@ 3.4 Odor discrimination with a spiking neural network model
The dataset presented above allows us to estimate the sensor responses for different sets of odorants. We now proceed further down our processing pipeline by modeling the neural network stage. To do so, we train a spiking neural network (SNN) to classify odors based on the electrode responses. The SNN architecture used for this task consists of three fully connected layers. The number of input neurons is determined by the number of different ORs used as the sensing front-end, while each output neuron corresponds to a specific odor class. The neurons in the model follow the Leaky-Integrate-and-Fire (LIF) model, with the subthreshold dynamics described by the following equation:
@@
where $V_{mem}$ is the membrane potential of neuron $i$, $j$ is the presynaptic neuron index, $\sigma(t)$ is the spike at time $t$ from neuron $i$ to $j$, $\alpha$ is the membrane decay factor, $\Delta t$ is the time step and $C$ is the membrane capacitor for charge accumulation and leakage. When the membrane potential reaches the threshold value $V_{thr}$ the neuron emits one spike at time $t$.
@@
In this case study, the same odor sets (blue, orange, and green) used in the previous subsection were selected for classification (see Fig. 4). For each odorant, synthetic response traces were generated. The response values for each OR at each time step were organized into a response matrix, where the height of the matrix corresponds to the number of ORs and the width of the matrix represents the number of time steps. This time series was sequentially fed into the MLP SNN, column by column, with the analog value passed into the input-to-hidden linear layer, followed by an activation function which defines the behavior of membrane potential update. The following layer operates in the same manner. The spike outputs for each odor category were then summed across all time steps for each sample. The resulting test accuracy is shown in Fig. 6. As these results demonstrate, the SNN learns to classify the odor sets; however, as expected, the classification accuracy depends on the number of receptors and the chemical structure of the odors.
Figure 6: Test accuracy (median and quartiles) for classification of three different sets of odorants (outliers are marked in red squares), with varying numbers of receptors (3, 9, 15, and 21). The number of hidden layer neurons is fixed at 100. Training was performed 10 times for each set of odorants. All weights were quantized to 4 bits.
Note that the SNN classification results do not exactly follow the odor space analysis of Fig. 4; for example, the blue set is classified with lower accuracy, even though it has higher separability. We attribute this to our definition of separability of odors as the linear distance between odor responses. Therefore, a non-linear analysis like an SNN can exploit deeper underlying correlations, skewing the classification results across odor sets. These SNN results can be seen as baseline results, with more complex and biorealistic networks and algorithms required to tackle continual learning, drift compensation, and more complex odor combinations (see Discussion).
@@ 3.5 Electronic sensing pad and neuromorphic hardware implementation of the spiking neural network model
An illustration of the complete hardware architecture is shown in Fig. 7. The proposed system consists of an analog front-end designed to process chemosensory signals using an event-driven analog spiking neural network architecture. By bypassing analog-to-digital converters (ADCs) and by leveraging novel synthetic biological assemblies, this approach enables real-time, low-power operation suitable for always-on applications. The system is manufactured in the SkyWater 130 nm CMOS process and integrates: i) a bioelectro interface that couples lipid vesicles with synthetic olfactory receptors (OR) and microelectrode arrays, and ii) a low-noise analog amplifier for signal conditioning. iii) a neuromorphic spiking neural network that processes sensor responses in a low-power mixed-signal domain.
@@
Figure 7: Overview of the hardware implementation, including the layout and schematic view of individual blocks. The components are as follows: Low Noise Amplifier (LNA) (upper left), pads for the sensing front-end (upper left), and the neuron tile with synapses (upper and center right). The Ngspice simulation results are shown below: voltage update at node $V_{sin}$ (center left), membrane potential across the capacitor $C_{1}$ in the NEURON block updates for an equivalent three-bit resolution ranging from 100 mV to 700 mV in seven equal intervals named $w_{0},..w_{6}$ (bottom left), and membrane potential integration with low leakage (bottom right). The physical analog circuits support 4-bit analog weights ($W_{0,1,2,3}$), and results in seven distinguishable weight values ($w_{0},..w_{6}$).
The schematic overview of the neuron circuit is shown in Fig. 7 “NEURON” block. Three node voltages ($V_{dc}$, $V_{leak}$, and $V_{ref}$) are used to control the membrane potential $V_{mem}$, that is, the voltage across the capacitor $C_{1}$, influencing its charging speed, discharging speed, and refractory period, respectively. The behavior of $V_{mem}$ below the firing threshold is described by equation 1, in which the voltage at the node $V_{leak}$ corresponds to the decay of $\alpha$, and $V_{dc}$ controls a constant injection of current ($I_{dc}$ in equation 1). Due to this constant charging behavior, the model parameter $\alpha$ depends on $V_{mem}$ and is expressed as:
$\alpha=1-\frac{R_{leak}\cdot\Delta t}{V_{mem}}$ (2)
where $R_{leak}$ is the leakage rate (V/s), and $\Delta t$ is the chosen time step.
@@
The neuron receives input from the preceding layers of neurons through the synapse circuit, whose schematic is shown in Fig. 7 “SYNAPSE” block. The variables $W[0\sim 3]$ define the weight (current) for each bit, while $S[0\sim 3]$ act as switches to enable their respective contributions. The output current, $i_{out}$, represents the weighted spike contribution and is directly injected into $V_{mem}$ of the neuron circuit. The resulting charging behavior of $V_{mem}$ is illustrated in Fig. 7 (bottom). To balance the parameters between bias and weight matrix, the transistor sizes for the input $dc$ and the weighted spike are deliberately designed to be different.
The relationship between $V_{W[0\sim 3]}$ and the discharge rate $R_{V_{sin}}$ of the node $V_{sin}$ follows a simple subthreshold MOS transistor transfer characteristic, of which the equation is 3.
$R_{V_{sin}}=\frac{dV_{sin}}{dt}=\frac{I_{tot}}{C}=\frac{1}{C}\sum_{Branch\ i}\frac{I_{ds,0_{i}}W_{i}}{L_{i}}e^{\frac{{V_{W_{i}}}-V_{th,n}}{nU_{T}}}$ (3)
[…]
Subsequently, the node $V_{sin}$ is converted to a current $i_{out}$ by a P-type MOSFET $M_{0}$. Ultimately, $i_{out}$ charges the $V_{mem}$ of the neuron. This contribution corresponds to the second term, $\sum_{j\neq i}w_{i,j}\sigma_{j}(t)$, in Equation 1. It is important to note that the exponential characteristic of $M_{0}$ complicates the precise representation of the bit value.
@@ 3.6 Hybrid synthetic sensory neuron
We simulated the full signal processing pipeline from the ion dynamics on the microelectrode to the firing of VLSI spiking neurons. The microelectrode pad array is responsible for detecting low-amplitude bioelectronic signals in the tens of microvolt range (as shown in Fig. 8 top), which are then fed into a low-noise amplifier (LNA) for signal conditioning. The LNA enhances weak signals, providing an output swing of approximately 5-10 mV with a gain of 50.05 dB, ensuring reliable signal amplification while preserving fidelity (second plot in Fig. 8). This amplified signal is then processed by spiking neurons, where the modulated activity of current injection is provided by connecting the LNA output, after scaling, to the ($V_{dc}$) nodes, directly stimulating somatic circuits to generate spikes. The event-driven nature of this architecture enables efficient signal encoding, where spike generation is proportional to the detected signal strength, facilitating mean rate coding for robust event-based detection (three bottom plots in Figure 8). A comprehensive simulation captures the entire signal processing pipeline, where soma circuits are maintained near the threshold by balancing the contributions of injection and leakage currents ($V_{dc}$ and $V_{leak}$), allowing neurons to rapidly generate spikes upon detection of simulated odor signals on the pad.
@@
4 Conclusions & Outlook
This work presents a novel co-design methodology and a neuromorphic chemosensing system that aims at integrating biological olfactory receptors with analog neuromorphic circuits, mimicking natural olfactory processing. By combining biological odor receptors with neuromorphic hardware, we aim to achieve real-time, energy-efficient chemosensory detection. The use of weak-inversion transistors and event-driven spiking neural networks ensures on-demand processing, minimizing power consumption while maintaining high sensitivity to chemical stimuli.
As we have outlined, our approach is unique in that it combines the physical realization of sensory neurons, processing, and learning mechanisms, all in one device. To reach this goal, we build on recent advances in the use of machine learning methods in biology to enable the co-design with the established electrical engineering methods.
A key component is the bioelectronic interface, which bridges biological odor recognition with CMOS-based electronics. To enhance signal transduction, it will be necessary to characterize the synthetic cell-electrode interface and to optimize the electrode impedance for improved sensitivity and signal integrity.

### v2

@@ 3.2 Surface stabilization of Orco-OR complexes in synthetic cells
To address the challenge of the necessary co-design approach, we developed a quantitative model of the proposed assembly of lipid-membrane-stabilized Orco-OR ion channels on an electrode. In the model, a fixed number of ion channels generate a current that flows into the electrode impedance. The ion channel was modeled as a resistor with two conductance states, and the electrodes were considered as a parallel resistor-capacitor circuit (see Fig. 9B Pads & Equivalent Circuit). The cleft between the synthetic cell and the electrode acts as a resistive element and is relatively large due to the nanoscopic dimensions of the cleft (see Materials and Methods). Analysis of the model showed that the vesicle radius r is an important design parameter. The total charge in a synthetic cell scales with $r^{3}$, while the number of ion channels scales with $r^{2}$. Thus, we expect two regimes with different vesicle sizes: one regime dominated by the finite number of charges encapsulated in smaller vesicles, whereas signals from larger vesicles are sensitive to the density of inserted ion channels. In addition, the overall coverage of signal-generating synthetic cells depends on the ratio of electrode size to vesicle size. Vesicles sizes can be readily tuned between 100 nm and 100 $\mu$m, providing an opportunity to test the different regimes predicted by the model [106].
@@ 3.3 Selection methodology for odor receptors (OR) and synthetic dataset
To this end, we make use of a synthetic dataset based on measurements of odorant responses in Drosophila melanogaster, which we construct from the DoOR dataset.
We first need to restrict the number of ORs and odorants. The task is to find the minimal number of ORs we need to reconstitute in order to reliably differentiate between a given set of odors. This can be seen as a separation task in the space of responses of all receptors, see Fig. 4 (left). We achieve this by calculating the ‘odor separability’, i.e. the Euclidean distance between the response vectors to the odor, and iteratively excluding the OR which contributes least to separability; see Section 6.4 for a detailed description of the method.
Our analysis yields the odor separability for different sets of odorants vs. the number of enabled ORs, as shown in Fig. 4 (center). For illustration, we have chosen three sets of odors with differing degrees of relationship in terms of chemical structure (blue, orange, and green sets). As expected, all odor sets show better separability with larger number of different receptors; however, relative separability naturally depends on the combination of odors we are probing (note that results are normalized across odorant sets). Blue: three odorants from different functional groups (respectively: alcohols, esters, and lactones) show large relative separability. Orange: separability is reduced when considering less diverse functional groups (here: alcohols and aromatic alcohols). Green: odorants from a single functional group with similar OR responses are naturally less separable. Note that we control against the possibility of no ligand binding at all for a given odorant by including a ‘no odor’ category based on the spontaneous firing rate (SFR); see Section 6.4 in Materials and Methods.
Most importantly, our algorithm tells us exactly which ORs need to be reconstituted for optimal separability of odors and subsequent discrimination of odors by our sensing platform. As an example, Fig. 4 (right) shows the most relevant ORs for $N_{\text{OR}}=3$ for the aforementioned odor sets.
@@
Figure 4: Left: Odor separability, exemplified by three smells that occupy a space defined by the responses of different ORs. Our algorithm selects OR for highest separability in the OR/odor space, maximizing the distance between odor responses. Center: Normalized odor separability vs. number of Dm ORs for different sets of odorants and spontaneous firing rate (no odor, SFR). Right: ORs showing highest separability for each odor set for $N_{\text{OR}}=3$.
@@
We assume that each synthetic cell contains ORs of the same type, and model the response to the binding of an odor molecule as a stochastic process with $N_{\text{ORCO}}$ ion channels. If the ligand binds to a closed Orco channel, the channel opens, and current flows out of the vesicle until stochastic unbinding occurs. We thus define a Markov model with the opening probability of Orco $p_{\text{open}}$ and closing with $p_{\text{close}}$, see Fig. 5 (left). The open and closed states are converted to currents by comparison with the electrophysiological characterization of VUAA1-induced Orco currents; see Section 6.5 and Butterwick et al. 2018 [84], in particular Extended Data Fig. 2. Similarly, ion channel noise and low-pass filter properties of the vesicle membrane are modeled on the basis of electrophysiological characterization. For the electrode response, we assume that synthetic cells of different receptor types are physically separated, such that each electrode probes only one OR type. In addition, we parametrize the electrode model as described in Section 3.2. Analysis of the electrical circuit shows that the sensor response to an odor molecule is given by a low-pass filter of the summed Orco-OR currents in each neuron. From this procedure, we obtain a time series of voltages with sensor responses per odor–OR type combination, see Fig. 5. The model is flexible with respect to which odorants and ORs to simulate, supporting the full DoOR 2.0.1 dataset of 692 odors and 52 OR types. In order to increase realism w.r.t. actual measurement conditions, where the timing of odor presentation is not always known precisely, the presentation onset is chosen at random within the full simulation time window. To avoid accidental correlation of onset timing with odors, all simulations use different onsets for each recording, such that onsets are not odorant-specific.
Refer to caption Refer to caption
Figure 5: Left: Illustration of sensory neuron model. Each Orco-OR complex is modelled as an independent Markov process with two states. Right: Simulated electrode responses to three odorants. Each trace shows the response of one synthetic cell comprising 10 Orco-OR complexes of the same receptor type; different receptor types are color-coded. We show the 7 most relevant OR types, as determined by our OR reduction algorithm. Opacity of traces is scaled with maximum voltage for better readability. Odors are only presented during a portion of the overall time frame (dashed vertical line) – note that odors and onsets are however not correlated in the dataset (multiple presentations of same odor have different onsets).
@@ 3.4 Odor discrimination with a spiking neural network model
The dataset presented above allows us to estimate the sensor responses for different sets of odorants. We now proceed further down our processing pipeline by modeling the neural network stage. To do so, we train a spiking neural network (SNN) to classify odors based on the electrode responses. The SNN architecture used for this task consists of three fully connected layers (Fig. 6 left). The number of input neurons is determined by the number of different ORs used as the sensing front-end, while each output neuron corresponds to a specific odor class. The neurons in the model follow the Leaky-Integrate-and-Fire (LIF) model, with the subthreshold dynamics described by the following equation:
@@
where $V_{mem}$ is the membrane potential of neuron $i$, $j$ is the presynaptic neuron index, $\sigma(t)$ is the spike at time $t$ from neuron $i$ to $j$, $\alpha$ is the membrane decay factor, $w_{i,j}$ denotes the weight of the synapse connecting neuron $i$ to $j$, $\Delta t$ is the time step and $C$ is the membrane capacitor for charge accumulation and leakage. $I_{dc}$ describes the continuous current injected to the membrane from an analog current input. When the membrane potential reaches the threshold value $V_{thr}$ the neuron emits one spike at time $t$.
@@
In this case study, the same odor sets (blue, orange, and green) used in the previous subsection were selected for classification (see Fig. 4). For each odorant, synthetic response traces were generated. The response values for each OR at each time step were organized into a response matrix, where the height of the matrix corresponds to the number of ORs and the width of the matrix represents the number of time steps. This time series was sequentially fed into the multi-layer SNN, column by column, with the analog value passed into the input-to-hidden linear layer, followed by an activation function which defines the behavior of membrane potential update. The following layer operates in the same manner. The spike outputs for each odor category were then summed across all time steps for each sample. The resulting test accuracy is shown in Fig. 6 (right). As these results demonstrate, the SNN learns to classify the odor sets; however, as expected, the classification accuracy depends on the number of receptors and the chemical structure of the odors.
Figure 6: An illustration of the SNN with test accuracy (median and quartiles) for classification of three different sets of odorants (outliers are marked in red squares), with varying numbers of receptors (3, 9, 15, and 21). Each OR group contains 100 samples per odorant, including the “SFR” category, resulting in a total of 400 samples. The dataset is split into 70% training, 15% validation and 15% testing. Test accuracy reported corresponds to the epoch with the highest validation accuracy during the 15-epoch training. The number of hidden layer neurons is fixed at 50. Training was performed 10 times for each set of odorants. All weights are quantized to 4 bits.
Note that the SNN classification results do not exactly follow the odor space analysis of Fig. 4; for example, the blue set is classified with lower accuracy, even though it has higher separability. We attribute this to our definition of separability of odors as the linear distance between odor responses. Therefore, a non-linear analysis like an SNN can exploit deeper underlying correlations, skewing the classification results across odor sets. These SNN results can be seen as baseline results, with more complex and biorealistic networks and algorithms required to tackle continual learning, drift compensation, and more complex odor combinations (see Discussion & Outlook).
@@ 3.5 Electronic sensing pad and neuromorphic hardware implementation of the spiking neural network model
An illustration of the complete hardware architecture is shown in Fig. 7. The proposed system consists of an analog front-end designed to process chemosensory signals using an event-driven analog spiking neural network architecture. By bypassing analog-to-digital converters (ADCs) and by leveraging novel synthetic biological assemblies, this approach enables real-time, low-power operation suitable for always-on applications. The system is designed in the SkyWater 130 nm CMOS process and integrates: i) a bioelectro interface that couples lipid vesicles with synthetic olfactory receptors (OR) and microelectrode arrays, and ii) a low-noise analog amplifier for signal conditioning. iii) a neuromorphic spiking neural network that processes sensor responses in a low-power mixed-signal domain.
@@
Figure 7: Overview of the hardware implementation, including the layout and schematic view of individual blocks. The components are as follows: Low Noise Amplifier (LNA) (A), pads for the sensing front-end and its equivalent circuit schematic (B), and the neuron tile with synapses (D and E). The Ngspice simulation results are shown below: voltage update at node $V_{sin}$ (G), membrane potential across the capacitor $C_{1}$ (equivalence of membrane capacitor) in the NEURON block (E) updates for an equivalent three-bit resolution ranging from 100 mV to 700 mV in seven equal intervals named $w_{0},..w_{6}$ (H), and membrane potential integration with low leakage (I). The physical analog circuits support 4-bit binary weight combinations through $S_{0,1,2,3}$ (each switch can be either "1" or "0"), and the binary combinations result in seven distinguishable weight values $w_{0},..w_{6}$.
@@ 3.6 Hybrid synthetic sensory neuron
We simulated the full signal processing pipeline from the ion dynamics on the microelectrode to the firing of VLSI spiking neurons. The microelectrode pad array is responsible for detecting low-amplitude bioelectronic signals in the tens of microvolt range (as shown in Fig. 8 top), which are then fed into a low-noise amplifier (LNA) for signal conditioning. The LNA enhances weak signals, providing an output swing of approximately 5-10 mV with a gain of 50.05 dB, ensuring reliable signal amplification while preserving fidelity (second plot in Fig. 8). This amplified signal is then processed by spiking neurons, where the modulated activity of current injection is provided by connecting the LNA output, after scaling, to the ($V_{dc}$) nodes, directly stimulating somatic circuits to generate spikes. The event-driven nature of this architecture enables efficient signal encoding, where spike generation is proportional to the detected signal strength, facilitating mean rate coding for robust event-based detection (three bottom plots in Fig. 8). A comprehensive simulation captures the entire signal processing pipeline, where soma circuits are maintained near the threshold by balancing the contributions of injection and leakage currents ($V_{dc}$ and $V_{leak}$), allowing neurons to rapidly generate spikes upon detection of simulated odor signals on the pad.
@@
4 Discussion & Outlook
A key component of our platform is the bioelectronic interface, which bridges biological odor recognition with CMOS-based electronics. To enhance signal transduction, it will be necessary to characterize the synthetic cell-electrode interface and optimize the electrode impedance for improved sensitivity and signal integrity. We are relatively confident to be able to obtain a high electrical signal fidelity, as systematic engineering of electrodes for biosignal detection is well established, for example, by polymer layers [109]. A larger conceptual problem is the limitation to sensing of biocompatible molecules, e.g. avoiding substances that might disintegrate the synthetic cells, and in general the limitation in the lifetime of biological receptors outside of living systems. After all, there will be good reasons why living cells invest in turnover (disintegration and synthesis) of membrane proteins with typical half-lives of days [110]. We therefore believe that disposable sensor elements are a realistic first implementation of our design. Due to the low power consumption of neuromorphic systems, enabling compact design, many uses of disposable sensors come to mind in the domains of environmental sampling, safety, and food processing [17]. Moreover, by integrating several chips into the platform, each microelectrode operating in its own bath and targeting a single OR type, the system could be scaled into a multi-microelectrode architecture. This arrangement would allow simultaneous screening of multiple ORs while still allowing incremental testing as needed.